OpenStudy (anonymous):

Find the value of a and b for the curve x^2y+ay^2=b if the point (1,1) is on its graph and the tangent line at (1,1) has the equation 4x+3y=7

OpenStudy (stacey):

By plugging (1, 1) into the equation, we get 1+a=b. Do you know how to find the tangent line, or even how to find the derivative of the equation x^2y+ay^2=b?

OpenStudy (anonymous):

Hint: you can set the derivative of your function at \((1,1)\) equal to the slope of the tangent line.

OpenStudy (saranya):

This can be solved because of the included bit about the tangent line, because now you have enough information to create 2 simultaneous equations. You have, x^(2y) + ay^2 = b, and the other is found by finding its derivative. 2y*x^(2y - 1) + 2ay * dy/dx = 0 And we know what x, y, and dy/dx are. x is 1 and y is 1 from the coordinate (1,1) dy/dx can be found with the tangent equation, by expressing it in slope-intercept form. 4x + 3y = 7 3y = 7 - 4x y = 7/3 - (4/3)x So, dy/dx is 4/3. Now subsitute x, y and dy/dx into each equation, and simplify it. x^(2y) + ay^2 = b (1)^(2(1)) + a(1)^2 = b 1 + a = b 2y*x^(2y - 1) + 2ay * dy/dx = 0 2(1)*(1)^(2(1) - 1) + 2a(1)(4/3) = 0 2 + 8a/3 = 0 You can now solve for a. 2 + 8a/3 = 0 8a/3 = -2 8a = -6 a = -3/4 And now back substituting for a to find b. 1 + a = b 1 + (-3/4) = b b = 1/4 Therefore, a = -3/4 b = 1/4